# Certification of high stellar ranks of quantum states of light with a pair of click detectors

### v3

@@ Abstract
Stellar rank of quantum states of light quantifies the amount of non-Gaussian resources required for their generation. One popular and practical approach to certification of stellar rank is based on measurement of click statistics with an array of binary detectors that can only distinguish the presence and absence of photons. Specifically, it was shown that measurements with an array of $m+1$ detectors allow one to certify stellar rank of approximate Fock state $|m\rangle$, even when the state is subjected to losses or certain noise. Here we address the question how many click detectors are in principle required to certify high stellar ranks. We show that two click detectors arranged in a Hanbury Brown-Twiss setup suffice. Interestingly, detection of higher stellar rank is greatly facilitated by making the total detection efficiency of the detectors sufficiently low but well calibrated. Losses affect the response of the detection scheme in a way that can be exploited to certify stellar rank of high Fock states. We explicitly construct the corresponding stellar rank witnesses and discuss dependence of the stellar rank thresholds on parameters of the considered setup. Our results reveal that it is possible to certify high stellar ranks even with a minimalistic scheme that provides only very coarse grained information about the photon number statistics of the characterized state.
@@ I Introduction
Generation of highly non-classical states of light such as Fock states [1, 2, 3, 4], various cat-like states [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], approximate Gottesman-Kitaev-Preskill states [16, 17], states exhibiting non-linear squeezing [18, 19], or other interesting states [20, 21, 22] requires non-Gaussian operations [23, 24, 25] in addition to Gaussian squeezing, coherent displacements, passive linear optics, and homodyne detection [26, 27]. The non-Gaussian resources required for preparation of a given quantum state of light can be quantified by its stellar rank [28, 29]. Finite superpositions of Fock states possess stellar rank equal to the number of the highest Fock state in the superposition. Stellar rank is invariant under Gaussian unitary operations and Gaussian states have stellar rank equal to zero. There exist states with infinite stellar rank such as superpositions of coherent states. The concept of stellar rank is equivalent to the genuine $n$-photon quantum non-Gaussianity introduced in Ref. [30].
@@
In practical terms, the stellar rank quantifies the number of elementary photon additions [31, 32, 33, 34, 35] or photon subtractions [5, 6, 8] that have to be performed to generate the state, following e.g. the protocols proposed in Refs. [36, 37]. Equivalently, the stellar rank defines the minimum number of photons that have to be detected when the state is prepared by photon counting measurements on a part of multimode Gaussian entangled state [38]. The simplest yet illustrative example is the conditional preparation of Fock state $|n\rangle$ by detection of $|n\rangle$ photons in the idler mode of two-mode squeezed vacuum state [1, 2, 3, 4].
@@
Stellar rank can be certified by suitable stellar rank witnesses [30, 29, 39]. Various witnesses of stellar rank have been proposed in the literature and tested experimentally [29, 39, 30, 40, 41, 43, 42, 44, 45]. The stellar rank witnesses represent generalization and extension of witnesses of quantum non-Gaussianity [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Those latter witnesses can certify that the state is not a Gaussian state or a mixture of Gaussian states and thus possesses non-zero stellar rank.
@@
Quantum states of light are often detected with binary click detectors that can only distinguish the presence or absence of photons. To enable approximate photon counting one can utilize spatial or temporal multiplexing [60, 61, 62, 63, 64, 68, 65, 67, 66, 69, 70, 74, 71, 72, 73, 30, 75, 76, 77, 79, 78, 80] and split the input signal onto an array of $M$ detectors and measure the probabilities that a particular set of $k$ detectors click. In Ref. [30], stellar-rank witnesses based on linear combinations of clicks of $m$ detectors and clicks of $m+1$ detectors for array of $m+1$ click detectors were constructed and utilized to certify the stellar rank (or equivalently the genuine $m$-photon quantum non-Gaussianity) of multipohoton Fock states.
@@
Certification of stellar rank $m$ of Fock state $|m\rangle$ by coincidence measurements with $M>m$ binary click detectors is a natural choice since such detection reasonably well approximates the ideal photon counting. One can nevertheless wonder whether it would be possible to certify stellar rank of Fock states even with simpler experimental configuration and smaller number of binary detectors. Here we answer this question in affirmative and we show that it suffices to employ a scheme with two binary detectors. Interestingly, detection of high stellar rank is greatly facilitated by making the total detection efficiency sufficiently small which can be ensured by inserting additional losses in front of the detectors. It has been proposed in the past that photon number distribution of a quantum state can be reconstructed from measurements with a single binary click detector preceded with tunable losses [81, 82]. By contrast, here the losses are fixed and a single measurement configuration is utilized to certify the stellar rank. While the losses are helpful, they are not strictly necessary and we show that an unbalanced detection scheme can certify stellar rank higher than $1$ even with perfect detectors. However, even in this case losses can make the detection of higher stellar ranks more experimentally feasible.
@@
Recently, quantum non-Gaussianity criteria that account for losses of click detectors were presented in Ref. [83] and utilized to certify quantum non-Gaussianity of photons emitted from quantum dot. Inclusion of detection efficiency into the description of the detector makes the resulting quantum non-Gaussianity criteria more efficient since the losses that belong to the detection are not attributed to the state which is characterized [83]. Here we focus on different aspect of lossy detection, namely the ability to certify high stellar rank with a small number of click detectors. Admittedly, the considered setup can certify high stellar ranks of a narrow subset of states only. Nevertheless, we believe that it is of fundamental interest to investigate what are the minimum experimental resources required for certification of higher stellar ranks, and the setup considered in the present work is remarkably simple.
@@
The rest of the paper is organized as follows. In Section II we review some basic definitions and facts concerning the stellar rank and its certification. We also specify and discuss a property of quantum measurement diagonal in Fock basis that makes it useful for certification of higher stellar ranks. In Section III we describe the considered experimental scheme. In Section IV we present the numerically derived stellar rank thresholds for stellar rank witness based on the probability that one detector clicks and the other does not click, and we discuss dependence of the thresholds on the setup parameters. In Section V we extend our analysis to more refined class of witnesses based on combination of probabilities of clicks of one detector and both detectors. Finally, Section VI contains a brief summary and conclusions.
@@ II Stellar rank
where $\hat{U}_{G}$ denotes an arbitrary Gaussan unitary operation, $|n\rangle$ stands for Fock states and $c_{m}\neq 0$.
In practice the generated quantum states are affected by various imperfections, noise and decoherence. It is therefore important to generalize the concept of stellar rank to arbitrary mixed states described by density matrix $\hat{\rho}$. A mixed state $\hat{\rho}$ has a stellar rank $m$ if it can be expressed as a statistical mixture of pure states with stellar rank at most $m$ but cannot be expressed as a mixture of pure states with stellar rank lower than $m$ only [28, 29]. More formally and rigorously, consider all possible pure state decompositions $\hat{\rho}=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, $p_{i}\geq 0$, and denote by $m_{i}$ the stellar rank of $|\psi_{i}\rangle$. The stellar rank $m_{\hat{\rho}}$ of mixed state $\hat{\rho}$ can be defined as [29]
@@
In experiments, we are interested in characterization and certification of stellar rank of the generated states. As already noted in the Introduction, various witnesses of stellar rank have been proposed and utilized in the literature. The simplest witnesses are based on fidelity $F$ of a given state with some pure state with stellar rank $m$ such as the Fock state $|m\rangle$ [29, 40, 34, 41]. More generally, on can consider a linear witness specified by some Hermitian operator $\hat{W}$ [29, 39, 30] and determine stellar rank thresholds $W_{m}$ such that the stellar rank $m$ is certified if
@@
Consider now a measurement scheme described by POVM elements $\hat{\Pi}_{k}$ that are diagonal in Fock basis. Let $p_{k}(n)=\langle n|\hat{\Pi}_{k}|n\rangle$ denote the probability of measurement outcome $k$ for input Fock state $|n\rangle$. Let us consider a witness formed by a specific POVM element, $\hat{W}=\hat{\Pi}_{k}$. A sufficient condition for this witness to certify the stellar rank of Fock state $|m\rangle$ is that
@@
Nonzero $\Delta$ ensures that the measured probability $p_{k}=\mathrm{Tr}[\hat{\Pi}_{k}\hat{\rho}]$ provides a lower bound on the fidelity $F_{m}$ with Fock state $|m\rangle$,
$F_{m}\geq 1-\frac{p_{k}(m)-p_{k}}{\Delta}.$ (6)
In order to certify the stellar rank $m$ based on the measured probability $p_{k}$ one can determine the corresponding stellar rank thresholds for the witness $\hat{W}=\hat{\Pi}_{k}$ as [29, 39]
$W_{m}=\sup_{|\phi_{m}\rangle,\,\hat{U}_{G}}\langle\phi_{m}|\hat{U}_{G}\hat{W}\hat{U}_{G}^{\dagger}|\phi_{m}\rangle,$ (7)
where the supremum is taken over all unitary Gaussian operations $\hat{U}_{G}$ and over all finite superpositions of Fock states $|\phi_{m}\rangle=\sum_{n=0}^{m-1}c_{n}|n\rangle$. Optimization over the pure states $|\phi_{m}\rangle$ leads to the maximum eigenvalue of the operator $\hat{U}_{G}\hat{\Pi}_{k}\hat{U}_{G}^{\dagger}$ projected onto the subspace $\mathcal{H}_{m-1}$ spanned by the Fock states up to $m-1$. Let us define the corresponding projector
$\hat{Z}_{m-1}=\sum_{n=0}^{m-1}|n\rangle\langle n|.$ (8)
We have [39, 84]
$W_{m}=\sup_{\hat{U}_{G}}\left\{\max\mathrm{eig}\left[\hat{Z}_{m-1}\hat{U}_{G}\hat{\Pi}_{k}\hat{U}_{G}^{\dagger}\hat{Z}_{m-1}\right]\right\}.$ (9)
The remaining optimization over all single-mode Gaussian unitaries can be performed numerically.
Figure 1: Proposed measurement scheme (a). The input signal is attenuated by attenuator Att and split at a beam splitter BS to be detected by two binary click detectors D1 and D2. Example of tunable attenuator for input linearly polarized light (b). The attenuator is formed by a sequence of half-wave plate HWP and polarizing beam splitter PBS.
@@ III Measurement scheme
After these general considerations, let us focus on the experimental scheme depicted in Fig. 1. The input signal is first attenuated by an attenuator with intensity transmittance $\eta_{A}$ and then it is split into two parts at a beam splitter BS with transmittance $T$ and reflectance $1-T$. The transmitted and reflected modes are detected with binary click detectors D1 and D2 that distinguish the presence and absence of photons. The detection efficiency $\eta_{D}$ of the detectors can be incorporated into the overall loss factor $\eta=\eta_{A}\eta_{D}$. The POVM element associated with a click of the binary on/off detector with total detection efficiency $\eta$ reads
@@
The probability $R_{1}(n)$ that only detector D1 in Fig. 1(a) clicks can be expressed as a difference of probability that D2 does not click irrespective of the response of D1, and the probability that both D1 and D2 do not click. For input Fock state $|n\rangle$ we obtain
@@
In order to get some insight into the mechanism which shifts the maximum detection probabilities to higher Fock states it is instructive to compare results for Fock states $|1\rangle$ and $|2\rangle$. To simplify the discussion it is convenient to attribute all losses to the attenuator Att. For input Fock state $|1\rangle$ the probability that D1 clicks reads $R_{1}(1)=\eta T$, which is the product of probabilities that a single photon passes through the attenuator and reaches the correct output port with detector D1. For the input Fock state $|2\rangle$ both photons will pass through the attenuator with probability $\eta^{2}$ and with probability $2\eta(1-\eta)$ only one of the photons passes through the attenuator. Consequently, the probability $R_{1}(2)$ can be expressed as a sum of two contributions,
@@
Observe that for $\eta<\frac{1}{2}$ the attenuated two-photon Fock state contains higher single-photon contribution than the attenuated single-photon state, $2\eta(1-\eta)>\eta$. More generally, by comparinng the single-click probabilities $R_{1}(1)$ and $R_{1}(2)$ we find that the peak of the function $R_{1}(n)$ will be at $n\geq 2$ when
@@
Figure 2: Probabilities $R_{1}(n)$ of click of detector D1 and simultaneous no-click of D2 for input Fock states $|n\rangle$ are plotted for balanced detection scheme with $T=0.5$ and four different total detection efficiencies $\eta=1$ (a), $\eta=0.6$ (b), $\eta=0.4$ (c) and $\eta=0.2$ (d).
The dependence of $R_{1}(n)$ on $T$ is plotted in Fig. 3 for fixed $\eta=2/3$, which illustrates the role of the beam splitter transmittance. We can see that as $T$ increases the peak of $R_{1}(n)$ shifts to higher photon numbers $n$.
Figure 3: Probability $R_{1}(n)$ of click of detector D1 and simultaneous no-click of D2 for input Fock state $|n\rangle$ is plotted as function of $T$ for $\eta=2/3$ and $n=1$ (blue solid line), $n=2$ (red dashed line), and $n=3$ (green dot-dashed line).

### v4

# arXiv:2506.18391v4[quant-ph] 02 Mar 2026

@@ Abstract
Stellar rank of quantum state of light quantifes the amount of non-Gaussian resources required for its generation. One popular and practical approach to certification of stellar rank is based on measurement of click statistics with an array of binary detectors that can only distinguish the presence and absence of photons. Specifically, it was shown that measurements with an array of $m+1$ detectors allow one to certify stellar rank $m$ of approximate Fock state $|m\rangle$, even when the state is subjected to losses or certain noise. Here we address the question as to how many click detectors are in principle required to certify stellar ranks higher than one. We show that two click detectors arranged in a Hanbury Brown-Twiss setup suffice. Interestingly, detection of stellar ranks higher than one is greatly facilitated by making the total detection efficiency of the detectors sufficiently low but well calibrated. Losses affect the response of the detection scheme in a way that can be exploited to certify stellar rank of higher Fock states. We explicitly construct the corresponding stellar rank witnesses and discuss dependence of the stellar rank thresholds on parameters of the considered setup. Our results reveal that it is possible to certify stellar ranks higher than $1$ even with a minimalistic scheme that provides only very coarse-grained information about the photon number statistics of the characterized state.
@@ I Introduction
Generation of highly non-classical states of light such as Fock states [1, 2, 3, 4], various catlike states [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], approximate Gottesman-Kitaev-Preskill states [16, 17], states exhibiting nonlinear squeezing [18, 19], or other interesting states [20, 21, 22] requires non-Gaussian operations [23, 24, 25] in addition to Gaussian squeezing, coherent displacements, passive linear optics, and homodyne detection [26, 27]. The non-Gaussian resources required for preparation of a given quantum state of light can be quantified by its stellar rank [28, 29]. Finite superpositions of Fock states possess stellar rank equal to the number of the highest Fock state in the superposition. Stellar rank is invariant under Gaussian unitary operations and Gaussian states have stellar rank equal to zero. There exist states with infinite stellar rank such as superpositions of coherent states. The concept of stellar rank is equivalent to the genuine $n$-photon quantum non-Gaussianity introduced in Ref. [30].
@@
In practical terms, the stellar rank quantifies the number of elementary photon additions [31, 32, 33, 34, 35] or photon subtractions [5, 6, 8] that have to be performed to generate the state, following, e.g., the protocols proposed in Refs. [36, 37]. Equivalently, the stellar rank defines the minimum number of photons that have to be detected when the state is prepared by photon counting measurements on a part of multimode Gaussian entangled state [38]. The simplest yet illustrative example is the conditional preparation of Fock state $|n\rangle$ by detection of $n$ photons in the idler mode of two-mode squeezed vacuum state [1, 2, 3, 4].
@@
Stellar rank can be certified by suitable stellar rank witnesses [30, 29, 39]. Various witnesses of stellar rank have been proposed in the literature and tested experimentally [29, 39, 30, 40, 41, 43, 42, 44, 45]. The stellar rank witnesses represent generalization and extension of witnesses of quantum non-Gaussianity [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Those latter witnesses can certify that the state is not a Gaussian state or a mixture of Gaussian states and thus possesses nonzero stellar rank.
@@
Quantum states of light are often detected with binary click detectors that can only distinguish the presence or absence of photons. To enable approximate photon counting one can utilize spatial or temporal multiplexing [60, 61, 62, 63, 64, 68, 65, 67, 66, 69, 70, 74, 71, 72, 73, 30, 75, 76, 77, 79, 78, 80] and split the input signal onto an array of $M$ detectors and measure the probabilities that a particular set of $k$ detectors click. In Ref. [30], stellar rank witnesses based on linear combinations of clicks of $m$ detectors and clicks of $m+1$ detectors for array of $m+1$ click detectors were constructed and utilized to certify the stellar rank $m$ (or equivalently the genuine $m$-photon quantum non-Gaussianity) of multipohoton Fock states.
@@
Certification of stellar rank $m$ of Fock state $|m\rangle$ by coincidence measurements with $M>m$ binary click detectors is a natural choice since such detection reasonably well approximates the ideal photon counting. One can nevertheless wonder whether it would be possible to certify stellar rank of Fock states even with simpler experimental configuration and smaller number of binary detectors. Here we answer this question in affirmative and we show that it suffices to employ a scheme with two binary detectors. Interestingly, detection of stellar ranks higher than one is greatly facilitated by making the total detection efficiency sufficiently small which can be ensured by inserting additional losses in front of the detectors. It has been proposed in the past that photon number distribution of a quantum state can be reconstructed from measurements with a single binary click detector preceded with tunable losses [81, 82]. By contrast, here the losses are fixed and a single measurement configuration is utilized to certify the stellar rank. While the losses are helpful, they are not strictly necessary and we show that an unbalanced detection scheme can certify stellar rank higher than one even with perfect detectors. However, even in this case losses can make the detection of higher stellar ranks more experimentally feasible.
@@
Recently, quantum non-Gaussianity criteria (i.e., criteria for certification of stellar rank one) that account for losses of click detectors were presented in Ref. [83] and utilized to certify quantum non-Gaussianity of photons emitted from quantum dot. Inclusion of detection efficiency into the description of the detector makes the resulting quantum non-Gaussianity criteria more efficient since the losses that belong to the detection are not attributed to the state which is characterized [83]. Here we focus on a different aspect of lossy detection, namely, the ability to certify stellar ranks higher than one with a minimum number of click detectors. Admittedly, the considered setup can certify high stellar ranks of a narrow subset of states only. Nevertheless, we believe that it is of fundamental interest to investigate what are the minimum experimental resources required for certification of higher stellar ranks, and the setup considered in this work is remarkably simple.
@@
The rest of the paper is organized as follows. In Sec. II we review some basic definitions and facts concerning the stellar rank and its certification. We also specify and discuss a property of quantum measurement diagonal in Fock basis that makes it useful for certification of higher stellar ranks. In Sec. III we describe the considered measurement scheme. In Sec. IV we present the numerically derived stellar rank thresholds for stellar rank witness based on the probability that one detector clicks and the other does not click, and we discuss dependence of the thresholds on the setup parameters. In Section V we extend our analysis to more refined class of witnesses based on combination of probabilities of clicks of one detector and both detectors. Various experimental factors that can influence the ability to certify higher stellar ranks are discussed in Sec. VI. Finally, Sec. VII contains a brief summary and conclusions.
@@ II Stellar rank
where $\hat{U}_{G}$ denotes an arbitrary Gaussan unitary operation, $|n\rangle$ stands for Fock states, and $c_{m}\neq 0$.
In practice the generated quantum states are affected by various imperfections, noise, and decoherence. It is therefore important to generalize the concept of stellar rank to arbitrary mixed states described by density matrix $\hat{\rho}$. A mixed state $\hat{\rho}$ has a stellar rank $m$ if it can be expressed as a statistical mixture of pure states with stellar rank at most $m$ but cannot be expressed as a mixture of pure states with stellar rank lower than $m$ only [28, 29]. More formally and rigorously, consider all possible pure-state decompositions $\hat{\rho}=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, $p_{i}\geq 0$, and denote by $m_{i}$ the stellar rank of $|\psi_{i}\rangle$. The stellar rank $m_{\hat{\rho}}$ of mixed state $\hat{\rho}$ can be defined as [29]
@@
In experiments, we are interested in characterization and certification of stellar rank of the generated states. As already noted in the Introduction, various witnesses of stellar rank have been proposed and utilized in the literature. The simplest witnesses are based on fidelity $F$ of a given state with some pure state with stellar rank $m$ such as the Fock state $|m\rangle$ [29, 40, 34, 41]. More generally, one can consider a linear witness specified by some Hermitian operator $\hat{W}$ [29, 39, 30] and determine stellar rank thresholds $W_{m}$ such that the stellar rank $m$ is certified if
@@
The stellar rank witnesses are conceptually similar to entanglement witnesses [84, 85] since in both cases the goal is to prove that a given quantum state $\hat{\rho}$ does not belong to some convex set of states. In the present case these convex sets $\mathcal{S}_{m}$ are formed by all pure states (1) and their convex mixtures. Therefore, the witness threshold $W_{m}$ can be determined by maximization of $\mathrm{Tr}[\hat{W}\hat{\rho}$] over $\mathcal{S}_{m-1}$. Thanks to the linearity of the witness, it actually suffices to perform the maximization over all pure states (1) [29, 39].
More explicitly, we can express $W_{m}$ as [29, 39]
$\color[rgb]{0,0,0}W_{m}=\sup_{|\phi_{m-1}\rangle,\,\hat{U}_{G}}\langle\phi_{m-1}|\hat{U}_{G}\hat{W}\hat{U}_{G}^{\dagger}|\phi_{m-1}\rangle,$} (4)
where the supremum is taken over all unitary Gaussian operations $\hat{U}_{G}$ and over all finite superpositions of Fock states $|\phi_{m-1}\rangle=\sum_{n=0}^{m-1}c_{n}|n\rangle$. Optimization over the pure states $|\phi_{m-1}\rangle$ leads to the maximum eigenvalue of the operator $\hat{U}_{G}\hat{W}\hat{U}_{G}^{\dagger}$ projected onto the subspace $\mathcal{H}_{m-1}$ spanned by the Fock states up to m1m-1. Let us define the corresponding projector
$\hat{Z}_{m-1}=\sum_{n=0}^{m-1}|n\rangle\langle n|.$ (5)
We have [39, 86]
$\color[rgb]{0,0,0}W_{m}=\sup_{\hat{U}_{G}}\left\{\max\mathrm{eig}\left[\hat{Z}_{m-1}\hat{U}_{G}\hat{W}\hat{U}_{G}^{\dagger}\hat{Z}_{m-1}\right]\right\}.$} (6)
The remaining optimization over all single-mode Gaussian unitaries can be performed numerically, as discussed in more detail in Sec. IV.
Consider now a measurement scheme described by POVM elements $\hat{\Pi}_{k}$ that are diagonal in Fock basis. Let $p_{k}(n)=\langle n|\hat{\Pi}_{k}|n\rangle$ denote the probability of measurement outcome $k$ for input Fock state $|n\rangle$. Let us consider a witness formed by a specific POVM element $\hat{W}=\hat{\Pi}_{k}$. A sufficient condition for this witness to certify the stellar rank $m$ of Fock state $|m\rangle$ is that
@@
Nonzero $\Delta$ ensures that the measured probability $p_{k}=\mathrm{Tr}[\hat{\Pi}_{k}\hat{\rho}$] provides a lower bound on the fidelity $F_{m}$ with Fock state $|m\rangle$:
$F_{m}=\langle m|\hat{\rho}|m\rangle\geq 1-\frac{p_{k}(m)-p_{k}}{\Delta}.$ (9)
As shown in Refs. [46, 29] fidelities with Fock state represent witnesses of quantum non-Gaussianity and stellar rank, associated with witness operator $\hat{W}=|m\rangle\langle m|$. In particular, for any Fock state $|m\rangle$ it is possible to determine a threshold $F_{m,\mathrm{th}}$ such that the state has a stellar rank $m$ if $F_{m}>F_{m,\mathrm{th}}$, and the fidelity thresholds are strictly smaller than $1$. For any state, $p_{k}\geq p_{k}(m)F_{m}$ holds. This, together with Eq. (9), implies that the lower bound on fidelity $F_{m}$ will never be lower than
$F_{m,\mathrm{min}}=1-\frac{p_{k}(m)}{\Delta}(1-F_{m}),$ (10)
where $F_{m}$ is the true value of the fidelity. Stellar rank $m$ will thus be certified for any state that satisfies $F_{m,\mathrm{min}}>F_{m,\mathrm{th}}$, which can be equivalently expressed as
$F_{m}>1-\frac{\Delta}{p_{k}(m)}(1-F_{m,\mathrm{th}}).$ (11)
Note that, depending on the exact shape of the photon number distribution of state $\hat{\rho}$, certification of stellar rank $m$ may be possible even for fidelities $F_{m}$ smaller than the bound in Eq. (11).
Figure 1: Proposed measurement scheme (a). The input signal is attenuated by attenuator (Att) and split at a beam splitter (BS) to be detected by two binary click detectors D1 and D2. Example of tunable attenuator for input linearly polarized light (b). The attenuator is formed by a sequence of half-wave plate (HWP) and polarizing beam splitter (PBS).
@@ III Measurement scheme
After these general considerations, let us focus on the experimental scheme depicted in Fig. 1. The input signal is first attenuated by an attenuator with intensity transmittance $\eta_{A}$ and then it is split into two parts at a beam splitter (BS) with transmittance $T$ and reflectance 1T1-T. The transmitted and reflected modes are detected with binary click detectors D1 and D2 that distinguish the presence and absence of photons. The detection efficiency $\eta_{D}$ of the detectors can be incorporated into the overall loss factor $\eta=\eta_{A}\eta_{D}$. The POVM element associated with a click of the binary detector with total detection efficiency $\eta$ reads as
@@
The probability $R_{1}$ that only detector D1 in Fig. 1(a) clicks can be expressed as a difference of probability that D2 does not click irrespective of the response of D1, and the probability that both D1 and D2 do not click. For input Fock state $|n\rangle$ we obtain
@@
In order to get some insight into the mechanism which shifts the maximum detection probabilities to higher Fock states it is instructive to compare results for Fock states $|1\rangle$ and $|2\rangle$. To simplify the discussion it is convenient to attribute all losses to the attenuator Att. For input Fock state $|1\rangle$ the probability that D1 clicks reads as $R_{1}(1)=\eta T$, which is the product of probabilities that a single photon passes through the attenuator and reaches the correct output port with detector D1. For the input Fock state $|2\rangle$ both photons will pass through the attenuator with probability $\eta^{2}$ and with probability $2\eta(1-\eta)$ only one of the photons passes through the attenuator. Consequently, the probability $R_{1}(2)$ can be expressed as a sum of two contributions,
@@
Observe that for $\eta<\frac{1}{2}$ the attenuated two-photon Fock state contains higher single-photon contribution than the attenuated single-photon state since $2\eta(1-\eta)>\eta$. More generally, by comparinng the single-click probabilities $R_{1}(1)$ and $R_{1}(2)$ we find that the peak of the function $R_{1}(n)$ will be at $n\geq 2$ when
@@
Figure 2: Probabilities $R_{1}(n)$ of click of detector D1 and simultaneous no-click of D2 for input Fock states $|n\rangle$ are plotted for balanced detection scheme with $T=0.5$ and four different total detection efficiencies $\eta=1$ (a), $\eta=0.6$ (b), $\eta=0.4$ (c), and $\eta=0.2$ (d).
The dependence of $R_{1}(n)$ on $T$ is plotted in Fig. 3 for fixed $\eta=\frac{2}{3}$, which illustrates the role of the beam-splitter transmittance. We can see that as $T$ increases the peak of $R_{1}(n)$ shifts to higher photon numbers $n$.
Figure 3: Probability $R_{1}(n)$ of click of detector D1 and simultaneous no-click of D2 for input Fock state $|n\rangle$ is plotted as function of $T$ for $\eta=\frac{2}{3}$ and $n=1$ (blue solid line), $n=2$ (red dashed line), and $n=3$ (green dotted-dashed line).
In order to obtain additional insight into the dependence of the shape of $R_{1}(n)$ on $\eta$ and $T$ we shall treat for a while $n$ as continuous parameter and seek $n$ that maximizes $R_{1}(n)$ by solving the extremal equation $\frac{dR_{1}(n)}{dn}=0$. This yields
$n_{\mathrm{max}}=\frac{\ln\displaystyle{\frac{\ln[1-\eta(1-T)]}{\ln(1-\eta)}}}{\ln(1-\eta)-\ln[1-\eta(1-T)]}.$ (17)
In practice, we should consider the $n_{\mathrm{max}}$ rounded up and down to find the integer number of photons which maximizes $R_{1}(n)$. Formula (17) simplifies in the limit of $\eta\ll 1$,
$n_{\mathrm{max}}\approx\frac{\ln(1-T)^{-1}}{\eta T}.$ (18)
Formulas (17) and (18) approximately quantify the maximum stellar rank that can be certified based on the knowledge of the probability $R_{1}$. In accordance with the results plotted in Figs. 2 and 3 these formulas indicate that the maximum certifiable stellar rank increases with increasing losses and also increases with increasing $T$.
The mechanism by which the losses help to increase the certifiable stellar rank is not limited to the specific detection scheme in Fig. 1 but is rather general. To illustrate this, we consider now a projection on Fock state $|k\rangle$ preceded by losses. The probability of detection of $k$ photons for input $n$-photon Fock state $|n\rangle$ reads as
$P_{k}(n)={n\choose k}\eta^{k}(1-\eta)^{n-k},\qquad n\geq k,$ (19)
and the probability vanishes for $n<k$. Specifically for $k=1$ we have
$P_{1}(n)=n\eta(1-\eta)^{n-1},$ (20)
and the maximum of $P_{1}(n)$ can be found analytically. After some algebra we get
$n_{\mathrm{max}}=-\frac{1}{\ln(1-\eta)}.$ (21)
In the limit of low $\eta$ we obtain $n_{\mathrm{max}}\approx 1/\eta$. For $k>1$ a similar approximate asymptotic formula can be derived, $n_{\mathrm{max}}\approx k/\eta$. The scaling $n_{\mathrm{max}}\propto 1/\eta$ thus emerges as rather universal.